# Predicting the suitability of photocatalysts for water splitting using Koopmans spectral functionals: The case of TiO2 polymorphs

### v1

@@ I Introduction
To better understand the desirable properties for a candidate photocatalyst, let us first briefly revise the PWS process. Water splitting is, of course, not a spontaneous reaction: it requires an energy barrier to be overcome. One of the ways to do this is via solar irradiation. Upon light absorption, the catalyst plays a crucial role in a three-step process: 1. charge carrier generation, in which electrons are promoted from the valence to the conduction band (creating a hole); 2. charge separation of the electron-hole pair and migration to the surface of a catalyst; and 3. participation in redox reactions. The first two steps are especially dependent on the structural and electronic properties of the photocatalytic material. The overall process is given by:
[…]
@@
The water-splitting reaction involves the oxidation of water and reduction of hydrogennder standard conditions the value of the redox potentials — as referenced to the normal hydrogen electrode — are 1.23 V and 0 V respectively [8]. Given these potentials, the band gap of a suitable catalyst needs to be at least 1.23 eV; otherwise the electrons will not have enough energy to start the reaction. In practice, at least 1.6 to 1.8 eV are required, as a certain amount of excess energy (i.e. the so-called “overpotential”) is needed to overcome kinetic barriers and induce the hydrogen and oxygen evolution reactions on the surface of the electrode [9, 10, 11]. The band gap should not be too high either, since that would reduce the amount of visible light that the photocatalyst can harness.
@@ II Theoretical framework
As mentioned, the conduction and valence band position with respect to the water redox levels is one of the most important features of a promising photocatalyst. The positions of the band edges depend on — among other factors — the crystal structure, the chemical environment, and the surface structure of the material [35, 36, 37]. Broadly speaking, “band alignment” refers to measuring the relative alignment of the energy bands of two materials at an interface. Often this interface is between two solid materials (such as semiconductors) but it can also refer to a junction between a semiconductor and fluid (such as the case of water splitting via photocatalytic surface). There are several methods for calculating the alignment of bands at an interface, which can be broadly categorized based on the choice of reference level, or, relatedly, what system is modeled (e.g. the bulk alone vs. the explicit interface). The band alignment can be estimated from the bulk structure alone via branch point energies (BPEs) [38]. BPEs are defined as the energy points where bands change character from being predominantly donor-like to acceptor-like. It has been shown that the electronegativity and variation in character of interface-induced gap states are determined by the nearest band edge, which can in turn be used to estimate the band offset [39]. One clear advantage of this method is that it only requires calculations of the bulk, drastically reducing its computational cost, but the method cannot be applied to materials for which the BPEs lie in the conduction band region [40]. At the other end of the spectrum, the hetero-interface can be explicitly modeled. For hetero-interfaces without lattice mismatch, this approach is straightforward because it only requires aligning the valence and conduction bands of semiconductors against the calculated average potentials in the plane parallel to the interfaces [41]. In the case of lattice mismatch, potential deformation of a core state needs to be taken into account [42, 43]. Finally, an intermediate approach is to measure the band offsets with respect to a reference level. This reference level can be the vacuum level, in which one obtains the ionization potential (IP) and electron affinity (EA) of each material [44]. These properties are intrinsically surface properties and must be obtained by considering a slab model. Alternatively, an impurity can be used as a reference level, such as hydrogen [45] or transition metal impurities [46], where band alignment relies on the position of dangling bonds formed between the impurity and semiconductor. Other more approximate approaches exist, such as effective dipole moments, tight-binding schemes, and empirical schemes [47, 48]. All of the above methods are nicely summarized in Ref. 49. In this work, we use the method based on the IP and EA of the material, adapted for use with orbital-density-dependent functionals [50, 51, 52, 53].
@@ II.1 Band alignment with DFT
[…]
@@ II.2 Koopmans spectral functionals
So why are KS-DFT eigenvalues unreliable, and what can we do to improve them? To start to answer this question, we note that DFT Kohn-Sham eigenvalues do not match the corresponding energy difference one obtains from explicitly performing electron addition/removal (i.e. a $\Delta$SCF calculation). Contrast this with the exact one-body Green’s function — an object that describes one-particle excitations exactly — whose excitation energies (i.e. poles) are located precisely at points that correspond to addition/removal energies (as can be straightforwardly seen from its spectral representation).
@@
As this equation shows, Koopmans functionals take the form of a correction to DFT: The first term on the right-hand side, $E^{\mathrm{DFT}}$, is the energy of a (typically semi-local) “base” functional. The corrective terms then impose GPWL (equation 7) by removing, orbital-by-orbital, all non-linear dependence of the total energy on the orbital occupancies (the first term inside the parentheses) and replacing it with a term that is linear with respect to the orbital occupation $\emph{f}_{i}$ (the second term inside the parentheses) [56, 57, 58, 59]. Here $\eta_{i}$ is some constant. Depending on how one chooses $\eta_{i}$, different flavors of Koopmans can be defined, such as KI and KIPZ [56]. The Koopmans integral correction (KI) sets $\eta_{i}$ in order to guarantee that the orbital energies $\varepsilon_{i}$ are equal to the corresponding $\Delta$SCF total energy difference of the base functional. Explicitly, the KI functional is given by
@@
where $\rho_{i}$ is orbital density at filling $f_{i}$ and $n_{i}$ is the normalized orbital density i.e. $\rho_{i}(\mathbf{r})=f_{i}n_{i}(\mathbf{r})$. The parameters $\alpha_{i}$ are screening parameters, which account for the fact that the derivation of the KI correction only accounts for the explicit dependence of the DFT energy on the orbital occupancies in going from equation 8 to 9. By scaling the strength of the orbital correction via these screening parameters, we account for orbital relaxation post hoc. Crucially, these screening parameters $\alpha_{i}$ are system-specific, and can be computed ab initio — i.e. they are not fitting parameters [60].
The second form of Koopmans functionals, KIPZ, takes the KI functional and adds one further term to the functional, a screened Perdew-Zunger (PZ) self-interaction correction term:
@@
which removes one-body self-interaction error and guarantees that the functional is exact for one-electron systems. Because evaluating the KIPZ functional is more expensive than the KI functional, it can be advantageous to evaluate the KIPZ Hamiltonian on the KI ground state. This is referred to as perturbative KIPZ (pKIPZ).
An important feature of Koopmans functionals is that they are not density functionals: the energy depends not only on the total density of the system but also on the densities of individual orbitals, making this an “orbital-density-dependent” functional theory. This comes with several important consequences. In the framework of DFT, the functional is invariant with respect to unitary rotations of the occupied manifold, and this means that the same set of orbitals minimize the total energy and diagonalize the Hamiltonian. However this is not the case for orbital-density dependent functionals, for which the “variational” orbitals minimize the energy while the “canonical” orbitals diagonalize the Hamiltonian constructed at the end of minimization procedure. These two sets of orbitals are related via a unitary transformation.
However it is important to note that emergence of localized variational orbitals is crucial because it allows Koopmans functionals to correctly treat bulk systems [57]. The Koopmans correction is applied to each variational orbital, while the canonical orbitals are interpreted as Dyson orbitals and their eigenvalues as quasiparticle energies, which are shifted by some amount depending on the size of the Koopmans corrections applied to each of their constituent variational orbitals. In the simplest case, where all the variational orbitals are symmetrically equivalent, the Koopmans correction results in a constant shift of the valence and conduction bands, opening the band gap. For systems with non-equivalent variational orbitals the Koopmans correction also affects inter- and intra-band distances, bandwidths, effective masses, etc..
@@
  • a Wannierization of the DFT Kohn-Sham states in order to obtain maximally-localized Wannier orbitals with which to initialize the variational orbitals
@@
The accuracy of Koopmans functionals has already been proven across a range of different materials [61, 63, 62, 64]. For more details about Koopmans functionals, we refer the reader to Ref. [65].
@@ II.3 Band alignment with Koopmans functionals
To calculate band alignment with Koopmans functionals, we must adapt the procedure described earlier in Section II.1. Here, we take advantage of a peculiar feature of the KI and pKIPZ functionals: for insulating systems where all the orbital occupancies are either 0 or 1, the ground state KI and pKIPZ energies and densities match those of the base functional. (This can be seen by setting $f_{i}$ to 0 or 1 in equation 9). We stress that this does not mean that the Koopmans correction has no effect, because the derivative of the correction is non-vanishing and thus the eigenvalues of the Koopmans functional will be different to those of the base functional. Nevertheless, this means that the ground-state density of Koopmans functionals match those of the base functional, and when performing band alignment the slab calculations only need to be performed at the level of DFT, because — being a ground-state property — the average electrostatic potential remains unchanged by the Koopmans correction [44, 66]. This substantially reduces the cost of this framework for computing band alignment.
Given this simplification, for Koopmans functionals equations 11 and 12 become:
@@
where $\Delta\varepsilon_{\mathrm{VBM}}$ and $\Delta\varepsilon_{\mathrm{VBM}}$ are the shifts in these band edges due to the Koopmans correction. Equivalently, one can calculate the electron affinity (EA) by simply subtracting the value of electronic gap calculated of Koopmans level from the IP:
@@ III Method
Koopmans functionals calculations were performed using Quantum ESPRESSO via the koopmans package [65, 75, 76]. The screening parameters were calculated via finite-differences, which necessitate the use of a supercell and image charge correction schemes to avoid spurious interactions between periodic images [77]. For these calculations, we used a $2\times 2\times 2$ supercell for rutile and anatase and $2\times 1\times 2$ supercell for brookite. To convert these $\Gamma$-only results to a well-sampled primitive cell the eigenvalues are unfolded to the equivalent primitive cell [63]. We used norm-conserving pseudopotentials from the SG15 library (version 1.0) [78]), and a wave-function energy cutoff of 80 Ry as recommended by force convergence data. All of the input and output files can be found on the Materials Cloud Archive [79].
@@
For the slab calculations, we used the thermodynamically most stable facets of $\mathrm{TiO_{2}}$ polymorphs: the (110) surface facet of rutile, (101) of anatase, and (210) of brookite. These surface orientations have been extensively studied in the literature [80, 81, 82, 83, 84, 85, 86, 87, 88] . The slab models were constructed using six and eight periodic units for rutile and anatase, and six periodic units for the larger brookite unit cell. The depth of the vacuum was chosen to be six times the height of the bulk primitive cell (ensuring correspondence in size between primitive and supercell), which corresponds somewhere between 20 and 30 Å depending on the polymorph. These slab sizes were chosen to ensure the macroscopic average potential was converged to within 0.01 eV. The slab calculations used a $3\times 7\times 1$ $k$-point mesh for rutile and anatase and $3\times 2\times 1$ for brookite. These DFT calculations were also performed using Quantum ESPRESSO. The geometries of the slab structures were optimized in order to obtain an accurate physical picture of the surface/vacuum interface.
@@ IV.1 Band structure
In Figure 3 we present the band structures of three polymorphs of TiO2 calculated via the KI functional, and compare them against those calculated with PBE (the underlying base functional). The band gaps of rutile and brookite TiO2 are direct ($\Gamma\rightarrow\Gamma$), while that of anatase is indirect ($Z\rightarrow\Gamma$) [89, 90, 91]. In all three cases, the KI correction shifts both the valence (downwards) and conduction bands (upwards), leading to the band gap widening relative to the PBE result. (The precise values of these shifts can be found in the Supplemental Material.) The resulting band gaps are presented in Table 1, alongside experimental values, as well as those calculated using hybrid functionals and GW. When comparing against experimental band gaps it is important to account for zero-point renormalization — by subtracting the ZPR (as calculated ab initio) from experimental values one obtains a value that can be fairly compared against the computational results obtained for pristine crystal geometries.
@@
G0W0@PBE11footnotemark: 1 $3.46$ $3.73$ $3.45$
G0W0@HSE0622footnotemark: 2 $4.73$- --
scQPGW@HSE0622footnotemark: 2 $5.18$- --
KI@PBE $3.29$ $3.80$ $4.03$
pKIPZ@PBE $3.38$ $3.90$ $4.14$
exp -- ZPR $3.34-3.41$ $3.34-3.44$-
exp333[93, 94, 95] $3.03-3.06$ $3.10-3.20$ $3.1-3.4$
$-0.337,$
ZPR444[96, 97, 98] $-0.349,$ $-0.238$-
$-0.314$
@@
The results with pKIPZ are very similar to that of KI (plots of pKIPZ@PBE band structures can be found in the Supplemental Material). Compared to the KI functional, pKIPZ reports slightly larger band gaps, especially in the case of brookite. Both KI and pKIPZ assign brookite the largest band gap of the three polymorphs, in contrast to HSE06 and G0W0@PBE which assign it the smallest. But without an accurate measurement of the ZPR in brookite and more precise experimental measurements of the band gap, it is difficult to determine which ordering is correct.

### v2

@@ I Introduction
To better understand the desirable properties for a candidate photocatalyst, let us first briefly revise the PWS process. The process is schematically presented in Figure 1.
Figure 1: Schematic illustration of photocatalytic water splitting. The conduction-band and valence-band regions are shown in green and red, respectively. When a photon (yellow line) with energy equal to or greater than the semiconductor band gap is absorbed, an electron is excited from the valence band to the conduction band, leaving a hole behind. The electron and hole subsequently participate in the reduction of hydrogen (top right) and oxidation of water (bottom right).
Water splitting is, of course, not a spontaneous reaction: it requires an energy barrier to be overcome. One of the ways to do this is via solar irradiation. Upon light absorption, the catalyst plays a crucial role in a three-step process: 1. charge carrier generation, in which electrons are promoted from the valence to the conduction band (creating a hole); 2. charge separation of the electron-hole pair and migration to the surface of a catalyst; and 3. participation in redox reactions. The first two steps are especially dependent on the structural and electronic properties of the photocatalytic material. The overall process is given by
$\begin{split}\mathrm{2H_{2}O(l)\longrightarrow O_{2}(g)+4H^{+}+4e^{-}\hskip 14.22636ptE_{ox}^{o}=1.23V}\end{split}$ (1)
@@
The water-splitting reaction involves the oxidation of water and reduction of hydrogen. Under standard conditions the value of the redox potentials — as referenced to the normal hydrogen electrode — are 1.23 V and 0 V respectively [96]. Given these potentials, the band gap of a suitable catalyst needs to be at least 1.23 eV; otherwise the electrons will not have enough energy to start the reaction. In practice, at least 1.6 to 1.8 eV are required, as a certain amount of excess energy (i.e. the so-called “overpotential”) is needed to overcome kinetic barriers and induce the hydrogen and oxygen evolution reactions on the surface of the electrode [53, 114, 87]. The band gap should not be too high either, since that would reduce the amount of visible light that the photocatalyst can harness.
@@ II Theoretical framework
As has already been mentioned, the valence and conduction band positions with respect to the water redox levels are some of the most important features of a promising photocatalyst. The positions of the band edges depend on — among other factors — the crystal structure, the chemical environment, and the surface structure of the material [88, 97, 93]. Broadly speaking, “band alignment” refers to measuring the relative alignment of the energy bands of two materials at an interface. Often this interface is between two solid materials (such as semiconductors) but it can also refer to a junction between a semiconductor and fluid (such as the case of water splitting via photocatalytic surface). There are several methods for calculating the alignment of bands at an interface, which can be broadly categorized based on the choice of reference level, or, relatedly, what system is modeled (e.g. the bulk alone vs. the explicit interface). The band alignment can be estimated from the bulk structure alone via branch point energies (BPEs) [105]. BPEs are defined as the energy points where bands change character from being predominantly donor-like to acceptor-like. It has been shown that the electronegativity and variation in character of interface-induced gap states are determined by the nearest band edge, which can in turn be used to estimate the band offset [82]. One clear advantage of this method is that it only requires calculations of the bulk, drastically reducing its computational cost, but the method cannot be applied to materials for which the BPEs lie in the conduction band region [11]. At the other end of the spectrum, the hetero-interface can be explicitly modeled. For hetero-interfaces without lattice mismatch, this approach is straightforward because it only requires aligning the valence and conduction bands of semiconductors against the calculated average potentials in the plane parallel to the interfaces [113]. In the case of lattice mismatch, potential deformation of a core state needs to be taken into account [68, 34]. Finally, an intermediate approach is to measure the band offsets with respect to a reference level. This reference level can be the vacuum level, in which one obtains the ionization potential (IP) and electron affinity (EA) of each material [102]. These properties are intrinsically surface properties and must be obtained by considering a slab model. Alternatively, an impurity can be used as a reference level, such as hydrogen [112] or transition metal impurities [64], where band alignment relies on the position of dangling bonds formed between the impurity and semiconductor. Other more approximate approaches exist, such as effective dipole moments, tight-binding schemes, and empirical schemes [20, 106]. All of the above methods are nicely summarized in Ref. 48. In this work, we use the method based on the IP and EA of the material, adapted for use with orbital-density-dependent functionals [55, 122, 100, 117].
@@ II.1 Band alignment with DFT
$\overline{\overline{V}}(z)=\frac{1}{L}\int_{z-\frac{L}{2}}^{z+\frac{L}{2}}\overline{V}(z^{\prime})dz^{\prime}$ (4)
@@ II.2 Koopmans spectral functionals
[…]
@@
$E_{\mathrm{Koopmans}}[\rho,{f_{i}}]=E^{\mathrm{DFT}}[\rho]+\sum_{i}\left(-\int_{0}^{f_{i}}\varepsilon_{i}(f)df+f_{i}\eta_{i}\right)$ (8)
As this equation shows, Koopmans functionals take the form of a correction to DFT: The first term on the right-hand side, $E^{\mathrm{DFT}}$, is the energy of a (typically semi-local) “base” functional. The corrective terms then impose GPWL (equation 7) by removing, orbital-by-orbital, all non-linear dependence of the total energy on the orbital occupancies (the first term inside the parentheses) and replacing it with a term that is linear with respect to the orbital occupation $\emph{f}_{i}$ (the second term inside the parentheses) [84, 85, 9, 10]. Here $\eta_{i}$ is some constant that will be defined shortly.
An important feature of Koopmans functionals is that they are not density functionals: the energy depends not only on the total density of the system but also on the densities of individual orbitals, making this an “orbital-density-dependent” functional theory. This comes with several important consequences. In the framework of DFT, the energy is invariant with respect to unitary rotations of the occupied manifold, and this means that the same set of orbitals minimize the total energy and diagonalize the Hamiltonian. This is not the case for orbital-density dependent functionals, for which the “variational” orbitals minimize the energy while the “canonical” orbitals diagonalize the Hamiltonian constructed at the end of minimization procedure. (Note that this does not imply that the functional is not variational. The loss of unitary invariance and the emergence of variational and canonical orbitals is a feature common to all orbital-density dependent functionals including e.g. PZ-SIC.)
@@
where $\rho_{i}$ is orbital density at filling $f_{i}$ and $n_{i}$ is the normalized orbital density i.e. $\rho_{i}(\mathbf{r})=f_{i}n_{i}(\mathbf{r})$. The parameters $\alpha_{i}$ are screening parameters, which account for the fact that the derivation of the KI correction only accounts for the explicit dependence of the DFT energy on the orbital occupancies in going from equation 8 to 9. By scaling the strength of the orbital correction via these screening parameters, we account for orbital relaxation post hoc. Crucially, these screening parameters $\alpha_{i}$ are system-specific, and can be computed ab initioi.e. they are not fitting parameters [18].
Curiously, for insulating systems where all the orbital occupancies are either 0 or 1, the ground state KI energies and densities match those of the base functional. (This can be seen by setting $f_{i}$ to 0 or 1 in equation 9, in which case the entire correction vanishes. (We stress that this does not mean that the Koopmans correction has no effect, because the derivative of the correction is non-vanishing and thus the eigenvalues of the Koopmans functional will be different to those of the base functional.) This property of the KI functional allows us to obtain the KI ground-state density directly from a single semi-local DFT calculation, after which only unitary rotations of the occupied variational orbitals are needed to locate the KI minimum. This greatly reduces the computational cost of these calculations.
The second form of Koopmans functionals, KIPZ, adds a screened Perdew-Zunger (PZ) self-interaction correction term to the functional:
@@
which removes one-body self-interaction error and guarantees that the functional is exact for one-electron systems. In contrast to KI, KIPZ does not share the same ground state density as the base functional and therefore mandates a full minimization of an orbital-density-dependent functional. Because of this, it can be advantageous to evaluate the KIPZ Hamiltonian on the KI ground state. This is referred to as perturbative KIPZ (pKIPZ).
@@
  2. 2.
    a Wannierization of the DFT Kohn–Sham states in order to obtain maximally localized Wannier orbitals with which to initialize the variational orbitals
  3. 3.
@@
There are a few important technical details regarding this procedure. First, while any unitary rotation of the occupied manifold leaves the DFT energy unchanged, ODD functionals assign different energies to different sets of orbitals with the same total density. As a consequence, the minimization cannot rely on the usual self-consistent diagonalization procedure. Instead, the variational orbitals are obtained by directly minimizing the ODD energy using two nested loops, following the strategy introduced in ensemble DFT [77]: an inner loop optimizes the orbital rotations at fixed density, while an outer loop allows the density itself to relax. This conjugate-gradient procedure yields the self-consistent minimum of the functional (see Supplemental Material S1 [1] for more detail). Additionally, note that orbital occupations are not treated as variational parameters in this procedure. Since all systems considered here have an electronic gap, the occupied orbitals always have $f_{i}=1$ and the empty orbitals $f_{i}=0$, and only the orbitals themselves are optimized.
Second, the emergence of variational orbitals — which are localized orbitals — is crucial because it allows Koopmans functionals to correctly treat bulk systems [85]. The Koopmans correction is applied to each variational orbital, while the canonical orbitals are interpreted as Dyson orbitals and their eigenvalues as quasiparticle energies, which are shifted by some amount depending on the size of the Koopmans corrections applied to each of their constituent variational orbitals. In the simplest case, where all the variational orbitals are symmetrically equivalent, the Koopmans correction results in a constant shift of the valence and conduction bands, opening the band gap. For systems with non-equivalent variational orbitals the Koopmans correction also affects inter- and intra-band distances, bandwidths, effective masses, etc.. This is discussed in more detail in Supplemental Material S2 [1].
Third, the computational cost of Koopmans functional calculations is dominated by the calculations to obtain the screening parameters. This procedure is detailed in Supplemental Material S3 [1], including a discussion of how these calculations scale. In brief, the calculations to obtain the screening parameters amount to computing the energies of various charged defects. In this context, image charge corrections are especially important (see Supplemental Material S4 [1]).
The accuracy of Koopmans functionals has already been demonstrated across a range of materials [19, 25, 17, 73]. For more details about Koopmans functionals, we refer the reader to Ref. [69].
@@ II.3 Band alignment with Koopmans functionals
To calculate band alignment with Koopmans functionals, we must adapt the procedure described earlier in Section II.1. Here, we take advantage of the aforementioned fact that ground state KI and pKIPZ energies and densities match those of the base functional. This means that when predicting band alignment with Koopmans functionals, the slab calculations only need to be performed at the level of DFT, because — being a ground-state property — the average electrostatic potential remains unchanged by the Koopmans correction [102, 99]. This substantially reduces the cost of this framework for computing band alignment.
However, it is important to note that this comes with a caveat. Since the ground-state energies of the base functional and the corrected ones (KI@PBE or pKIPZ@PBE) are identical, this implies that neither PBE, KI@PBE, nor pKIPZ@PBE can reproduce the experimental phase stability of the reported TiO2 polymorphs. That being said, the third Koopmans functional KIPZ does not share this property and therefore modifies both the ground-state energy and density, and could therefore be used to address relative phase stability more reliably. However, applying KIPZ in the context of this work would require performing all the slab calculations fully at the KIPZ level, significantly increasing the computational cost.
Proceeding with the KI and pKIPZ functionals, equations 11 and 12 become:
@@
where $\Delta\varepsilon_{\mathrm{VBM}}$ and $\Delta\varepsilon_{\mathrm{VBM}}$ are the shifts in the bulk band edges due to the Koopmans correction. Equivalently, one can calculate the electron affinity (EA) by simply subtracting the value of electronic gap calculated of Koopmans level from the IP:
@@ III Method
Koopmans functionals calculations were performed using Quantum ESPRESSO via the koopmans package [69, 39, 38]. The screening parameters were calculated via finite-differences, which necessitate the use of a supercell and image charge correction schemes to avoid spurious interactions between periodic images [76] (For more details, see Supplemental Material S4 [1]). For these calculations, we used a $2\times 2\times 2$ supercell for rutile and anatase and $2\times 1\times 2$ supercell for brookite. To convert these $\Gamma$-only results to a well-sampled primitive cell the eigenvalues are unfolded to the equivalent primitive cell [25]. We used norm-conserving pseudopotentials from the SG15 library (version 1.0) [98]), and a wave-function energy cutoff of 80 Ry as recommended by force convergence data. All of the input and output files can be found on the Materials Cloud Archive [103].
@@
For the slab calculations, we used the thermodynamically most stable facets of $\mathrm{TiO_{2}}$ polymorphs: the (110) surface facet of rutile, (101) of anatase, and (210) of brookite. These surface orientations have been extensively studied in the literature [108, 115, 75, 66, 28, 45, 2, 24, 42] . The slab models were constructed using six and eight periodic units for rutile and anatase, and six periodic units for the larger brookite unit cell. The depth of the vacuum was chosen to be six times the height of the bulk primitive cell (ensuring correspondence in size between primitive and supercell), which corresponds somewhere between 20 and 30 Å depending on the polymorph. These slab sizes were chosen to ensure the macroscopic average potential was converged to within 0.01 eV. The slab calculations used a $3\times 7\times 1$ $k$-point mesh for rutile and anatase and $3\times 2\times 1$ for brookite. These DFT calculations were also performed using Quantum ESPRESSO. The geometries of the slab structures were optimized using the PBE functional in order to obtain an accurate physical picture of the surface/vacuum interface.
@@ IV.1 Band structure
In Figure 4 we present the band structures of three polymorphs of TiO2 calculated via the KI functional, and compare them against those calculated with PBE (the underlying base functional). The band gaps of rutile and brookite TiO2 are direct ($\Gamma\rightarrow\Gamma$), while that of anatase is indirect ($Z\rightarrow\Gamma$) [27, 123, 72]. In all three cases, the KI correction opens the band gap, shifting the valence bands downwards and the conduction bands upwards. (The numerical shifts in the valence band maximum and conduction band minimum are provided in the Supplemental Material [1].) While the effect of the Koopmans correction may appear to be a rigid scissors shift, this is not the case; rather, it is an orbital-dependent correction whose effect on individual bands depend on the screening parameters and the orbital character. A full derivation of the KI potential and the conditions under which it reduces to a rigid shift is provided in Supplemental Material S2 [1].
The resulting band gaps of the titania polymorphs are presented in Table 1, alongside experimental values, as well as those calculated using hybrid functionals and GW. When comparing against experimental band gaps it is important to account for zero-point renormalization — by subtracting the ZPR (as calculated ab initio) from experimental values one obtains a value that can be fairly compared against the computational results obtained for pristine crystal geometries.
@@
G0W0@PBE11footnotemark: 1333[54] $3.66,3.46$ $4.03,3.73$ $3.45$
G0W0@HSE0622footnotemark: 2 $4.73$ - -
GW0@PBE33footnotemark: 3 $4.23$ $4.60$ -
QSGW444[14]22footnotemark: 2 $4.18,4.22$ - -
QSGW̃44footnotemark: 4555[21] $3.73,3.88$ - -
GW33footnotemark: 3 $4.84$ $5.28$ -
KI@PBE $4.04$ $4.51$ $4.63$
pKIPZ@PBE $4.19$ $4.64$ $4.78$
exp - ZPR 3.14.13.1-4.1 3.43.93.4-3.9 -
exp666[124, 119, 30, 107, 94, 90, 3, 33, 57, 58, 67, 95, 50] 2.83.82.8-3.8 3.23.73.2-3.7 3.13.53.1-3.5
0.337,-0.337,
ZPR777[80, 32, 118] 0.349,-0.349, 0.238-0.238 -
0.314-0.314
@@
[…]
Table 1 shows that the KI and KIPZ band gaps for rutile are in closer agreement with experiment than those for anatase and brookite, and that the differences among the three TiO2 polymorphs are more pronounced with KI and KIPZ than with HSE06 or $G_{0}W_{0}$. At present, no definitive explanation for this behavior can be offered. Given the wide spread in the reported experimental band gaps and the limited number of available data-points, drawing firm conclusions would be premature. There is no fundamental reason to expect differences between polymorphs to be inherently more pronounced when using Koopmans functionals than when using hybrid functionals or GW. Ongoing comprehensive benchmark studies are expected to clarify such trends more quantitatively.
Finally, without an accurate measurement of the ZPR in brookite and more precise experimental measurements of the band gap, it is difficult to determine which ordering is correct.